The project manager of a task force planning the construction of a domed stadium had hoped...

  1. The project manager of a task force planning the construction of a domed stadium had hoped to be able to complete construction prior to the next season. After reviewing construction duration estimates, it now appears that a certain amount of crashing will be needed to ensure project completion before the season opener. Given the following information, determine a minimum-cost crashing schedule that will shave five weeks off the project length.

Activity

Precedes

Normal Duration (Weeks)

CRASHING COSTS

First Week

CRASHING COSTS

Second Week

A

----

12

15000

$20000

B

A

14

10000

10000

C

----

10

5000

5000

D

CC

17

20000

21000

E

CC

18

16000

18000

F

CC

12

12000

15000

G

D

15

24000

24000

H

E

8

-----

------

I

F

7

30000

----

J

I

12

25000

25000

K

B

9

10000

10000

M

G

3

----

----

N

H

11

40000

-------

P

H, J

8

20000

20000

Assumptions:

  1. There is no minimum crash duration.
  2. Crashing cost varies by activity and by week i.e. Cash cost for 1st week, Crash cost for 2nd week.
  3. Activities without a crash cost cannot be crashed.
  4. No penalties and no incentives given.
Answer #1

Different paths and their duration is as below=

A-B-K=12+14+9=35

C-E-H-P=10+18+8+8=44

C-D-G-M=10+17+15+3=45

C-E-H-N=10+18+8+11=47

C-F-I-J-P=10+12+7+12+8=49

The below mentioned table has the details of crashing activities=

First of all we will crash activity C as it has least crashing cost

Path

0

1

2

3

4

5

A-B-K

35

C-E-H-P

44

43

C-D-G-M

45

44

C-E-H-N

47

46

C-F-I-J-P

49

478

Crashed activity

C

When it comes to second week, even then we will crash activity C due to least second week crash cost

Path

0

1

2

3

4

5

A-B-K

35

C-E-H-P

44

43

42

C-D-G-M

45

44

43

C-E-H-N

47

46

45

C-F-I-J-P

49

48

47

Crashed activity

C

C

Now in the remaining activities, F has the second least first week crashing cost so we will not crash Activity F

Path

0

1

2

3

4

5

A-B-K

35

C-E-H-P

44

43

42

41

C-D-G-M

45

44

43

42

C-E-H-N

47

46

45

44

C-F-I-J-P

49

48

47

46

Crashed activity

C

C

F

Again we will crash Activity f due to least second week crash cost after C activity

Path

0

1

2

3

4

5

A-B-K

35

C-E-H-P

44

43

42

42

42

C-D-G-M

45

44

43

43

C-E-H-N

47

46

45

45

45

C-F-I-J-P

49

48

47

46

45

Crashed activity

C

C

F

F

Final Table will be

Path

0

1

2

3

4

5

A-B-K

35

C-E-H-P

44

43

42

42

42

40

C-D-G-M

45

44

43

43

C-E-H-N

47

46

45

45

45

44

C-F-I-J-P

49

48

47

46

45

44

Crashed activity

C

C

F

F

E,P

Total cost of crashing=   5000+5000+12000+15000+3600=$73000